OpenStudy (anonymous):

Find tan(x) if sec^2(x) = 3/2? A. ±√(2)/2 B. ±1/4 C. ±√(6)/2 D. ±1/2

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sec^2x=\tan^2x+1 }\) \(\color{#000000 }{ \displaystyle \frac{3}{2}=\tan^2x+\frac{2}{2} }\) \(\color{#000000 }{ \displaystyle \tan^2x=\frac{1}{2} }\)

OpenStudy (solomonzelman):

The first line is a trigonometric identity. In second line I plug in 3/2 instead of sec²x, and I re-write 1 as 2/2. Then I subtract, 2/2 from both sides, and this way I am able to solve for tan²x.

OpenStudy (solomonzelman):

Now, you need to find tan(x), based on this.

OpenStudy (anonymous):

Thank you good sir! This is very thorough response :)

OpenStudy (solomonzelman):

You welcome!